# RainPro-8: An Efficient Deep Learning Model to Estimate Rainfall Probabilities Over 8 Hours

### v2

@@ Abstract
We present a deep learning model for high-resolution probabilistic precipitation forecasting over an 8-hour horizon in Europe, overcoming the limitations of radar-only deep learning models with short forecast lead times. Our model efficiently integrates multiple data sources - including radar, satellite, and physics-based numerical weather prediction (NWP) - while capturing long-range interactions, resulting in accurate forecasts with robust uncertainty quantification through consistent probabilistic maps. Featuring a compact architecture, it enables more efficient training and faster inference than existing models. Extensive experiments demonstrate that our model surpasses current operational NWP systems, extrapolation-based methods, and deep-learning nowcasting models, setting a new standard for high-resolution precipitation forecasting in Europe, ensuring a balance between accuracy, interpretability, and computational efficiency. Code is available at URL.
@@ 1 Introduction
Recent advances in artificial intelligence have generated significant interest in deep learning for weather forecasting (Rasp et al., 2024; An et al., 2024). Deep learning excels at handling complex, large-scale, high-dimensional data, making it well-suited for capturing intricate, nonlinear patterns in spatio-temporal systems (Manzhu Yu & Li, 2024). Although deep learning has achieved remarkable success in both nowcasting (Gao et al., 2024b; Gong et al., 2024) and medium-range forecasting (Lam et al., 2023; Price et al., 2024), significant challenges remain. Deep learning models for nowcasting are often limited to very short lead times (up to two hours). In contrast, medium-range models, which predict broader atmospheric dynamics for up to 10 days, typically operate at coarser resolutions and are influenced by precipitation-specific biases in the training datasets (Lavers et al., 2022). As a result, they struggle to capture small-scale precipitation features, like local showers, often leading to the exclusion of precipitation forecasts in medium-range models (Lam et al., 2023).
@@
  • Efficient architecture and training strategy for multi-source data of varying temporal or spatial resolution, consistent prediction of precipitation classes for all lead times at once.
  • Extensive empirical evaluation demonstrating that RainPro-8 outperforms existing operational methods by 65% and achieves significant gains over deep-learning nowcasting models; ablation studies and model attribution analysis of key design choices and model input.
  • Demonstration of RainPro’s versatility for radar-only 2-hour predictions on the SEVIR benchmark, achieving state-of-the-art performance compared to both deterministic and generative nowcasting models.
@@ 2 Related Work
Traditional methods for weather forecasting are mostly numerical weather prediction (NWP) models, which simulate atmospheric dynamics using mathematical equations, also as ensembles over multiple simulations (Toth & Kalnay, 1997). NWP models demand significant compute, especially for ensembles, which restricts their spatial and temporal resolution. Their spin-up time results in poor performance for short lead times (Ma et al., 2021). Extrapolation-based methods, e.g. PySTEPS (Pulkkinen et al., 2019), RainyMotion (Ayzel et al., 2019) are limited to reduced forecast lengths due to their assumptions on constant motion and intensity (van Nooten et al., 2023).
Deep learning formulates precipitation nowcasting as a spatio-temporal prediction (An et al., 2024). Approaches in this area include the use of convolutions in recurrent networks (Shi et al., 2015; 2017b; Wang et al., 2023b; Ma et al., 2024), the U-Net architecture (Ayzel et al., 2020; Fernández & Mehrkanoon, 2021; Zhang et al., 2023; Trebing et al., 2021), transformers (Gao et al., 2024b; Wu et al., 2024), diffusion models (Leinonen et al., 2023; Gao et al., 2024a; Yu et al., 2024; Gong et al., 2024), and adversarial training (Ravuri et al., 2021). Some work improves loss functions to tackle data imbalance, blurriness, and sparsity in precipitation nowcasting models (Xu et al., 2024a; Ko et al., 2022; Yan et al., 2024). Still, most efforts focus on short-term forecasts (1 to 3 hours) using only radar, where reduced lead time eliminates the need for multiple data sources, large contexts, or accounting for higher uncertainty associated with 8-hour predictions.
@@ 3.1 Probabilistic Precipitation Forecasting
$X=[R_{t}]^{0}_{t=-T_{\text{in}}+1}\in\mathbb{R}^{T_{\text{in}}\times H\times W}\qquad Y=[R_{t}]^{T_{\text{out}}}_{t=1}\in\mathbb{R}^{T_{\text{out}}\times H\times W},$ (1)
@@ 3.2 Ordinal Consistent Loss
Model output $P(R_{t}\geq\min(I_{c})|R_{t}\geq\min(I_{c-1}))$ for each intensity class $c$ guarantees monotonic probabilities $P_{t,c}$ that respect the inherent order of precipitation classes. For the lowest intensity class, the model outputs $P(R_{t}\geq\min(I_{1}))$. This approach, as in Fernandes & Cardoso (2018) for semantic segmentation, improves both interpretability and consistency in the generated forecasts. For any class, probability $P_{t,c}$ is the cumulative product of model outputs for all preceding classes:
@@ 3.3 Single-Pass Predictions
$\textrm{weights}_{\textrm{exp}}[t]=\exp(-\alpha\times t)\quad\textrm{weights}_{\textrm{norm}}[t]=\frac{\textrm{weights}_{\textrm{exp}}[t]}{\sum\textrm{weights}_{\textrm{exp}}[t]}\quad W_{t}=\frac{\textrm{weights}_{\textrm{norm}}[t]}{\overline{\textrm{weights}_{\textrm{norm}}}}\\$ (7)
@@ 3.4 Architecture
We design our architecture to forecast precipitation for a $512\times 512\textrm{km}^{2}$ patch. Assuming an average precipitation displacement rate of 1km per minute (Sønderby et al., 2020), we use 512km of spatial context on every side of the target patch ($\textrm{1536}^{2}\textrm{km}^{2}$) to provide our model with enough input context for all 8-hours of lead time. We use a U-Net (Ronneberger et al., 2015) with MaxViT (Tu et al., 2022) blocks to efficiently process multi-resolution data (Figure 1), similar to (Andrychowicz et al., 2023). Key differences include single-pass predictions without lead time conditioning (Section 3.3), early downsampling in the encoder, halving internal channels, and removing topographical embeddings, all contributing to a reduced parameter count of 36.7M from the original 227M.
@@
The architecture begins with data fusion on the encoder, using Space-to-Depth convolutions (Sunkara & Luo, 2023) and ResNet blocks (He et al., 2016). Input data is transformed from $BTCHW$ to $B(TC)HW$ and merged with sources at matching resolutions. To balance efficiency and performance, we only use the full input context of 512km at 8km and 16km inputs, and lower input context of 256km at 4km resolution. The encoder aggregates features for each resolution into a low-resolution representation while preserving skip connections for the U-Net decoder. To improve memory efficiency, we apply downsampling before ResNet blocks instead of after.
The low-resolution representation is processed by 12 MaxViT blocks. Based on Vision Transformer (Dosovitskiy et al., 2021), they combine local neighborhood and global gridded attention aggregate information across the full $\textrm{1536}^{2}\textrm{km}^{2}$ input patch. The decoder, using Transposed Convolutions (Long et al., 2015) and ResNet blocks, reconstructs high-resolution probability maps from the MaxViT output, while leveraging encoder skip connections. All forecast timesteps and intensity classes are generated by outputting $P(R_{t}\geq\min(I_{c})|R_{t}\geq\min(I_{c-1}))$ (Section 3.2) in a single forward pass with timestamps encoded into the channel dimension (Section 3.3). For geographical alignment, representations are padded and cropped at each resolution. E.g., 4km representations are padded equally on all sides before concatenating with 8km inputs, and MaxViT output is cropped to the target region with extra context for upsampling, later cropped to $\textrm{512}^{2}\textrm{km}^{2}$ for final output.
@@ 4 Experiments
RainPro-8 is implemented with PyTorch (Ansel et al., 2024) and PyTorch Lightning (Falcon & team, 2025), trained for 100k steps using a batch size of 16, validating every 2,000 steps. Training uses a static learning rate of 3e-4, AdamW optimizer with a weight decay of 0.1 and betas $(0.9,0.999)$, Exponential Moving Average (EMA) decay of 0.99975, dropout of 0.1, maximum stochastic depth (Huang et al., 2016) of 0.2, and lead time decay rate of 10. We use 256 channels throughout the entire network, totaling 36.7 million parameters. Any other network hyperparameters follow those of MetNet-3 (Andrychowicz et al., 2023). The model with the lowest validation loss is selected. Training is performed on an NVIDIA H100 SXM5 GPU and takes approx. 13 hours (cf. App. F).
@@
We use multiple data sources as input to our precipitation forecasting model, each matched to its supported spatial resolutions. RainViewer [^1] https://www.rainviewer.com/ radar composites serve as high-resolution ground truth at 4km and 8km resolutions to capture local detail and broader context. EUMETSAT [^2] https://user.eumetsat.int/data/satellites/meteosat-second-generation satellite imagery provides cloud-related features at 8km resolution, and NOAA’s GFS [^3] https://registry.opendata.aws/noaa-gfs-bdp-pds provides atmospheric variables and precipitation forecasts at 16km resolution. Topographical information from Copernicus DEM [^4] https://registry.opendata.aws/copernicus-dem is incorporated at 4km resolution. All data are resampled for spatial and temporal alignment and undergo normalization, clipping, and binning to handle varying scales and precipitation skewness. The data cover one year and over one million samples, with defined training, validation, and testing splits to ensure reliable model evaluation. Details on data, preprocessing, and samples in App. A.
@@
We use Critical Success Index (CSI) (Schaefer, 1990) at different thresholds, a standard accuracy metric in precipitation nowcasting (Andrychowicz et al., 2023; Gao et al., 2024b; Gong et al., 2024), and Continuous Ranked Probability Score (CRPS) (Hersbach, 2000), which assesses alignment of predicted probability distribution with observed values, rewarding sharp and reliable predictions. Fractions Skill Score (FSS) (Roberts & Lean, 2008) accounts for intensity shifts and offers tolerance to translation and deformation. Frequency Bias Index (FBI) (Termonia et al., 2018) quantifies over- and underforecasting, but not forecast quality. We also report Mean Absolute Error (MAE) and Mean Squared Error (MSE), but note that they are sensitive to high frequency of no-rain cases and thus less suitable for skewed distributions (Andrychowicz et al., 2023) (cf. App. B.1). Taking inspiration from the thresholding approach in MetNet (Espeholt et al., 2022), intensity-based metrics use the mean of the highest activated bucket of the cumulative output distribution. A bucket is activated if its predicted probability mass exceeds the corresponding threshold, where thresholds are computed for each bucket and lead time using the validation set. This approach effectively captures rare high-precipitation events, which tend to have lower predicted probability masses.
@@ 5 Conclusion & Future Work
6 Reproducibility Statement
The code will be made publicly available, along with instructions for accessing data via the official sources. The paper provides comprehensive details on the RainPro-8 framework, architecture, and evaluation, while the appendices serve as supporting material. Specifically, Section 3 describes the task definition, loss function, output generation, and model architecture. Section 4 summarizes the training setup and hyperparameters (with further details in Appendices E and F), data sources (including preprocessing and sample details in Appendices A and I), and baselines, metric definitions, and performance comparisons (expanded in Appendices B and G). Section 4 also presents visualizations (also in Appendix D), ablation studies, attribution analyses (also in Appendix C), and an evaluation of a RainPro-8 variant adapted for the benchmark task on a standard precipitation nowcasting benchmark (also in Appendix H).
@@ A.2 Data Preprocessing
We tackle challenges such as varying data ranges and skewed distribution of precipitation intensities using normalization, clipping, and binning. Min-max normalization is applied to all data sources, as many variables are non-normally distributed. Radar reflectivity values (dBZ) are clipped to a range of -1 to 64 to remove outliers. To address rainfall skewness, target radar maps are categorized into predefined rain intensity classes for multi-label classification. Using the Marshall-Palmer equation (Marshall & Palmer, 1948), dBZ values are converted to mm/h, with finer granularity for light rain ($\geq 0.1,\geq 0.2,\geq 0.4$) and broader ranges for heavier rainfall ($\textrm{up to}\geq 25.0\ \textrm{mm/h}$) (Appendix E). This approach captures subtle light rain variations while accounting for rare intense precipitation events, aligning with the original dBZ distribution.
@@ H.3 Evaluation
To assess the accuracy of precipitation nowcasting, we evaluate RainPro-2R against both deterministic and stochastic generative models. Deterministic baselines include SimVP (Gao et al., 2022) and Earthformer (Gao et al., 2024b), which employ a recurrent-free strategy to generate all frames simultaneously, as well as PhyDNet (Guen & Thome, 2020), MAU (Chang et al., 2021), and ConvGRU (Shi et al., 2017a), which generate frames sequentially using recurrent strategies. For state-of-the-art generative baselines, we include STRPM (Chang et al., 2022), MCVD(Voleti et al., 2022), and PreDiff (Gao et al., 2024a).

### v3

@@ Abstract
We present a deep learning model for high-resolution probabilistic precipitation forecasting over an 8-hour horizon in Europe, overcoming the limitations of radar-only deep learning models with short forecast lead times. Our model efficiently integrates multiple data sources - including radar, satellite, and physics-based numerical weather prediction (NWP) - while capturing long-range interactions, resulting in accurate forecasts with robust uncertainty quantification through consistent probabilistic maps. Featuring a compact architecture, it enables more efficient training and faster inference than existing models. Extensive experiments demonstrate that our model surpasses current operational NWP systems, extrapolation-based methods, and deep-learning nowcasting models, setting a new standard for high-resolution precipitation forecasting in Europe, ensuring a balance between accuracy, interpretability, and computational efficiency. Code is available at https://github.com/rafapablos/RainPro.
@@ 1 Introduction
Recent advances in artificial intelligence have generated significant interest in deep learning for weather forecasting (Rasp et al., 2024; An et al., 2024). Deep learning excels at handling complex, large-scale, high-dimensional data, making it well-suited for capturing intricate, nonlinear patterns in spatio-temporal systems (Manzhu Yu and Li, 2024). Although deep learning has achieved remarkable success in both nowcasting (Gao et al., 2024b; Gong et al., 2024) and medium-range forecasting (Lam et al., 2023; Price et al., 2024), significant challenges remain. Deep learning models for nowcasting are often limited to very short lead times (up to two hours). In contrast, medium-range models, which predict broader atmospheric dynamics for up to 10 days, typically operate at coarser resolutions and are influenced by precipitation-specific biases in the training datasets (Lavers et al., 2022). As a result, they struggle to capture small-scale precipitation features, like local showers, often leading to the exclusion of precipitation forecasts in medium-range models (Lam et al., 2023).
@@
  • We propose an efficient neural architecture and training strategy designed to integrate multi-source data with varying temporal and spatial resolutions. The model produces consistent probabilistic predictions across all forecast lead times in a single forward pass, enabling efficient and improved forecasts.
  • We conduct an extensive empirical evaluation demonstrating that RainPro-8 outperforms existing operational forecasting methods by 65% and achieves significant improvements over state-of-the-art deep learning nowcasting models. We further provide comprehensive ablation studies and model attribution analysis to quantify the impact of key design choices and input modalities.
  • We demonstrate the versatility of RainPro by adapting it to radar-only 2-hour precipitation prediction on the SEVIR benchmark, where it achieves state-of-the-art performance compared to both deterministic and generative nowcasting approaches.
@@ 2 Related Work
Traditional methods for weather forecasting are mostly numerical weather prediction (NWP) models, which simulate atmospheric dynamics using mathematical equations, also as ensembles over multiple simulations (Toth and Kalnay, 1997). NWP models demand significant compute, especially for ensembles, which restricts their spatial and temporal resolution. Their spin-up time results in poor performance for short lead times (Ma et al., 2021). Extrapolation-based methods, e.g. PySTEPS (Pulkkinen et al., 2019), RainyMotion (Ayzel et al., 2019) are limited to reduced forecast lengths due to their assumptions on constant motion and intensity (van Nooten et al., 2023).
Deep learning formulates precipitation nowcasting as a spatio-temporal prediction (An et al., 2024). Approaches in this area include the use of convolutions in recurrent networks (Shi et al., 2015; 2017b; Wang et al., 2023b; Ma et al., 2024), the U-Net architecture (Ayzel et al., 2020; Fernández and Mehrkanoon, 2021; Zhang et al., 2023; Trebing et al., 2021), transformers (Gao et al., 2024b; Wu et al., 2024), diffusion models (Leinonen et al., 2023; Gao et al., 2024a; Yu et al., 2024; Gong et al., 2024), and adversarial training (Ravuri et al., 2021). Some work improves loss functions to tackle data imbalance, blurriness, and sparsity in precipitation nowcasting models (Xu et al., 2024a; Ko et al., 2022; Yan et al., 2024). Still, most efforts focus on short-term forecasts (1 to 3 hours) using only radar, where reduced lead time eliminates the need for multiple data sources, large contexts, or accounting for higher uncertainty associated with 8-hour predictions.
@@ 3.1 Probabilistic Precipitation Forecasting
$X$ $=[R_{t}]^{0}_{t=-T_{\text{in}}+1}\in\mathbb{R}^{T_{\text{in}}\times H\times W},$ (1)
$Y$ $=[R_{t}]^{T_{\text{out}}}_{t=1}\in\mathbb{R}^{T_{\text{out}}\times H\times W},$
@@ 3.2 Ordinal Consistent Loss
Model output $P(R_{t}\geq\min(I_{c})|R_{t}\geq\min(I_{c-1}))$ for each intensity class $c$ guarantees monotonic probabilities $P_{t,c}$ that respect the inherent order of precipitation classes. For the lowest intensity class, the model outputs $P(R_{t}\geq\min(I_{1}))$. This approach, as in Fernandes and Cardoso (2018) for semantic segmentation, improves both interpretability and consistency in the generated forecasts. For any class, probability $P_{t,c}$ is the cumulative product of model outputs for all preceding classes:
@@ 3.3 Single-Pass Predictions
$\textrm{weights}_{\textrm{exp}}[t]=\exp(-\alpha\times t)$ (7)
$\textrm{weights}_{\textrm{norm}}[t]=\frac{\textrm{weights}_{\textrm{exp}}[t]}{\sum\textrm{weights}_{\textrm{exp}}[t]}$ (8)
$W_{t}=\frac{\textrm{weights}_{\textrm{norm}}[t]}{\overline{\textrm{weights}_{\textrm{norm}}}}\\$ (9)
@@ 3.4 Architecture
We design our architecture to forecast precipitation for a $512\times 512\textrm{km}^{2}$ patch. Assuming an average precipitation displacement rate of 1km per minute (Sønderby et al., 2020), we use 512km of spatial context on every side of the target patch ($\textrm{1536}^{2}\textrm{km}^{2}$) to provide our model with enough input context for all 8-hours of lead time. We use a U-Net (Ronneberger et al., 2015) with MaxViT (Tu et al., 2022) blocks to efficiently process multi-resolution data (Figure 1), similar to MetNet-3 (Andrychowicz et al., 2023). Key differences include single-pass predictions without lead time conditioning (Section 3.3), early downsampling in the encoder, halving internal channels, and removing topographical embeddings, all contributing to a reduced parameter count of 36.7M from the original 227M.
@@
The architecture begins with data fusion on the encoder, using Space-to-Depth convolutions (Sunkara and Luo, 2023) and ResNet blocks (He et al., 2016). Input data is transformed from $BTCHW$ to $B(TC)HW$ and merged with sources at matching resolutions. To balance efficiency and performance, we only use the full input context of 512km at 8km and 16km inputs, and lower input context of 256km at 4km resolution. The encoder aggregates features for each resolution into a low-resolution representation while preserving skip connections for the U-Net decoder. To improve memory efficiency, we apply downsampling before ResNet blocks instead of after.
The low-resolution representation is processed by 12 MaxViT blocks. Based on Vision Transformer (Dosovitskiy et al., 2021), they combine local neighborhood and global gridded attention aggregate information across the full $\textrm{1536}^{2}\textrm{km}^{2}$ input patch. The decoder, using Transposed Convolutions (Long et al., 2015) and ResNet blocks, reconstructs high-resolution probability maps from the MaxViT output, while leveraging encoder skip connections. All forecast timesteps and intensity classes are generated by outputting $P(R_{t}\geq\min(I_{c})|R_{t}\geq\min(I_{c-1}))$ (Section 3.2) in a single forward pass with timestamps encoded into the channel dimension (Section 3.3). For geographical alignment, representations are padded and cropped at each resolution. For example, 4km representations are padded equally on all sides before concatenating with 8km inputs, and MaxViT output is cropped to the target region with extra context for upsampling, later cropped to $\textrm{512}^{2}\textrm{km}^{2}$ for the final output.
@@ 4 Experiments
RainPro-8 is implemented with PyTorch (Ansel et al., 2024) and PyTorch Lightning (Falcon and team, 2025), trained for 100k steps using a batch size of 16, validating every 2,000 steps. Training uses a static learning rate of 3e-4, AdamW optimizer with a weight decay of 0.1 and betas $(0.9,0.999)$, Exponential Moving Average (EMA) decay of 0.99975, dropout of 0.1, maximum stochastic depth (Huang et al., 2016) of 0.2, and lead time decay rate of 10. We use 256 channels throughout the entire network, totaling 36.7 million parameters. Any other network hyperparameters follow those of MetNet-3 (Andrychowicz et al., 2023). The model with the lowest validation loss is selected. Training is performed on an NVIDIA H100 SXM5 GPU and requires approximately 13 hours (cf. App. F).
@@
We use multiple data sources as input to our precipitation forecasting model, each matched to its supported spatial resolutions. RainViewer [^1] https://www.rainviewer.com/ radar composites serve as high-resolution ground truth at 4km and 8km resolutions to capture local detail and broader context. Satellite imagery from EUMETSAT [^2] https://user.eumetsat.int/data/satellites/meteosat-second-generation provides cloud-related features at 8km resolution, and NOAA’s GFS [^3] https://registry.opendata.aws/noaa-gfs-bdp-pds provides atmospheric variables and precipitation forecasts at 16km resolution. Topographical information from Copernicus DEM [^4] https://registry.opendata.aws/copernicus-dem is incorporated at 4km resolution. To ensure consistency across these heterogeneous sources, all datasets are resampled to achieve spatial and temporal alignment. They further undergo normalization, clipping, and binning to handle varying scales and precipitation skewness. The data cover one year and over one million samples, with defined training, validation, and testing splits to ensure reliable model evaluation. Details on data, preprocessing, and samples in App. A.
@@
We use Critical Success Index (CSI) (Schaefer, 1990) at different thresholds, a standard accuracy metric in precipitation nowcasting (Andrychowicz et al., 2023; Gao et al., 2024b; Gong et al., 2024), and Continuous Ranked Probability Score (CRPS) (Hersbach, 2000), which assesses alignment of predicted probability distribution with observed values, rewarding sharp and reliable predictions. Fractions Skill Score (FSS) (Roberts and Lean, 2008) accounts for intensity shifts and offers tolerance to translation and deformation. Frequency Bias Index (FBI) (Termonia et al., 2018) quantifies over- and underforecasting, but not forecast quality. We also report Mean Absolute Error (MAE) and Mean Squared Error (MSE), but note that they are sensitive to high frequency of no-rain cases and thus less suitable for skewed distributions (Andrychowicz et al., 2023) (cf. App. B.1). Taking inspiration from the thresholding approach in MetNet (Espeholt et al., 2022), intensity-based metrics use the mean of the highest activated bucket of the cumulative output distribution. A bucket is activated if its predicted probability mass exceeds the corresponding threshold, where thresholds are computed for each bucket and lead time using the validation set. This approach effectively captures rare high-precipitation events, which tend to have lower predicted probability masses.
@@ 5 Conclusion & Future Work
Acknowledgments
This work is partly funded by the Innovation Fund Denmark (IFD) under File No. 2052-00064B. Computational resources were provided by Lambda and the GenomeDK cluster at Aarhus University.
Reproducibility Statement
The code is publicly available at https://github.com/rafapablos/RainPro, along with instructions for accessing data via the official sources. The paper provides comprehensive details on the RainPro-8 framework, architecture, and evaluation, while the appendices serve as supporting material. Specifically, Section 3 describes the task definition, loss function, output generation, and model architecture. Section 4 summarizes the training setup and hyperparameters (with further details in Appendices E and F), data sources (including preprocessing and sample details in Appendices A and I), and baselines, metric definitions, and performance comparisons (expanded in Appendices B and G). Section 4 also presents visualizations (also in Appendix D), ablation studies, attribution analyses (also in Appendix C), and an evaluation of a RainPro-8 variant adapted for the benchmark task on a standard precipitation nowcasting benchmark (also in Appendix H).
@@ A.2 Data Preprocessing
We tackle challenges such as varying data ranges and skewed distribution of precipitation intensities using normalization, clipping, and binning. Min-max normalization is applied to all data sources, as many variables are non-normally distributed. Radar reflectivity values (dBZ) are clipped to a range of -1 to 64 to remove outliers. To address rainfall skewness, target radar maps are categorized into predefined rain intensity classes for multi-label classification. Using the Marshall-Palmer equation (Marshall and Palmer, 1948), dBZ values are converted to mm/h, with finer granularity for light rain ($\geq 0.1,\geq 0.2,\geq 0.4$) and broader ranges for heavier rainfall ($\textrm{up to}\geq 25.0\ \textrm{mm/h}$) (Appendix E). This approach captures subtle light rain variations while accounting for rare intense precipitation events, aligning with the original dBZ distribution.
@@ H.3 Evaluation
To assess the accuracy of precipitation nowcasting, we evaluate RainPro-2R against both deterministic and stochastic generative models. Deterministic baselines include SimVP (Gao et al., 2022) and Earthformer (Gao et al., 2024b), which employ a recurrent-free strategy to generate all frames simultaneously, as well as PhyDNet (Guen and Thome, 2020), MAU (Chang et al., 2021), and ConvGRU (Shi et al., 2017a), which generate frames sequentially using recurrent strategies. For state-of-the-art generative baselines, we include STRPM (Chang et al., 2022), MCVD(Voleti et al., 2022), and PreDiff (Gao et al., 2024a).